# A ChatGPT-based approach for questions generation in higher education

### v1

@@ 1 Introduction
One of the most prominent advances in Artificial Intelligence recently is the development of large language models (LLMs), such as ChatGPT, BingChat, and Bard (developed by OpenAI, Microsoft, and Google, respectively). It can be said that large language models have been developing strongly in the past 2 years and creating a strong influence in the field of Generative AI (GenAI) Jahic2023.
LLMs are capable of solving a wide range of tasks, such as natural language understanding, text generation, and sentiment analysis in various domains chang2023survey. Since the breakthroughs of AI and LLM, education, as a crucial role in shaping society through almost every single individual, might receive significant benefits from these new LLM-based initiatives. AI is transforming education, bridging its gaps, and promoting a more inclusive and productive learning environment by customizing learning experiences, automating administrative tasks, and providing real-time feedback Kamalov.
Advancements in AI and LLMs have fueled the development of many educational technology innovations that aim to automate the often time-consuming and laborious tasks of generating and analyzing textual content, including generating test questions. For the application of LLM in education, this technology holds great potential in this field at all levels, especially university level Yan2023. Therefore, this study conducts a research to comprehensively understand the capabilities of LLMs, especially ChatGPT when supporting lecturer to build question banks for learning modules in university setting. During the literature review, we witnessed that although the LLMs can provide quite a lot of support in teaching activities Daoetal2023, the work of testing and evaluating learners, which takes up a lot of time and effort of lecturers, has not been studied in-depth by previous works. Therefore, in this paper, we decide to conduct research on the application of LLMs for generating question banks, with initial results at the Banking Academy of Vietnam. Specifically, the contributions of the paper are summarized as follows:
@@ 2 Related works
With the growing interest in LLM for education, many studies around the world focusing on this issue have been conducted and have produced some interesting results. A group of authors from the University of Minnesota Law School have conducted research on ChatGPT’s performance in answering a set of questions including 95 multiple-choice questions and 12 essay questions related to four undergraduate-level subjects in the law class. The results showed that ChatGPT studied and performed on the test at a level equivalent to a C+ grade for college students choi_hickman_monahan_schwarcz_2023.
Another study put ChatGPT to the test in completing the Dutch secondary school exam in the subject of English reading comprehension. This study concluded that ChatGPT achieved results with an average score of 7.3, comparable to the average score of all high school students in the country winter_2023.
In another research on evaluating the performance of LLMs when taking a high school level Biology test by the author namely Dao Xuan Quy, ChatGPT showed positive results with an accuracy of 71% for remembering level questions and 61.82% for understanding level questions. This result shows how effective ChatGPT is in capturing and clarifying concepts related to the required subject Dao2023.
The use of LLMs in supporting the construction of questions to test and evaluate students’ knowledge should be encouraged. While exam-style questions are an important instructional tool for several reasons, manually creating questions is a time-consuming process that necessitates expertise, experience, and resources. This, in turn, impedes and inhibits the implementation of instructional activities (e.g., offering practice questions) and educational initiatives (e.g., adaptive testing) that need a huge pool of questions. Therefore, automatic question generating (AQG) approaches based on AI were established for research in both developed and developing nations kurdi2019.
@@
Selecting the optimal LLM-based solution . Nowadays, many advanced large language models can be widely applied in the field of education such as ChatGPT, BingAI, Gemini, etc. Most prominently, ChatGPT is an advanced general AI model based on a Generative Pretrained Transformer (GPT). This feature refers to the LLM being pre-trained with the available dataset and can generategrammatically and contextually human-like responses according to the user’s request through natural language processing. The current most popular versions of ChatGPT are GPT-3.5 and GPT-4, with the upgraded version GPT-4 can be used with a fee, which according to OpenAI performs much better with a 10 times larger pre-training datasets capacity KALYAN2024100048.
The initial goal of this research is to evaluate various LLMs in order to find the one having performance and suitability for generating questions in the higher educational context. Each model has its own advantages and disadvantages, so our team has to compare consider which one can fulfill the both factors. As we prioritize the best ones meeting educational standards, ChatGPT Plus and Claude are two potential options. However, ChatGPT Plus has a relatively high fee (\$20/month) and Claude has a few limitations in its capabilities Lozi2023.
@@
  • ChatGPT 3.5 has been proven through scientific research to be able to produce high-quality text, similar to text written by humansRAY2023121.
  • ChatGPT 3.5 currently has 180 million users, with 100 million weekly active users and 1.6 billion website visit times meer2024. This number shows that ChatGPT 3.5 is sufficiently popular and trustworthy.
@@ 3.2 The investigation of appropriate Prompt patterns
Through research on the meaning of elements as well as the operation of available Prompt templates such as RTF (Role, Task, Format), RISE (Role, Input, Steps, Expectation), RTCF (Role, Task, Context, Format), RTCEF (Role, Task, Context, Example, Format). We have filtered, selected and separated into specific factors as shown in the Table 1 Cavoj2023.
@@ 3.5 The "Blind Test" method
To evaluate the quality of the questions more objectively, our research team conducted a survey using the “Blind test” method to collect answers from students to lecturers. Based on the well-known Turing Test, the "Blind test" aims to get empirical insight into people’s ability to discriminate between artificial and human content KOBIS2021106553.
@@ 3.6.3 Specific analysis
Our data shows that the 9 questions having the least percentage of students answering correctly have an average of only 32.5% answers correct. In particular, the survey showed that the multiple-choice section had the lowest rate of correct answers with only about 30%. This may stem from the fact that human lecturers often create short, concise multiple-choice questions with the goal of simply testing whether students have grasped the knowledge surrounding the concept being asked. In fact, LLM like ChatGPT has the ability to grasp the subject and create related questions to support learnersDao2023. Therefore, this tool is capable of creating MCQs properly along with the lecturers in an even faster way.
@@ D. Money loses value
Subjectively stated, a broad and general list of MCQs surveyed are less difficult to be guessed than real-life scenario exercises, or calculative ones that we had to give ChatGPT some examples to train it. Furthermore, these models can generate plausible distractors, which are incorrect answers that are designed to be similar to the correct answer, making it more challenging for humans to discern the difference. According to the study of Bitew et al. (2023), ChatGPT, when guided by question items and in-context examples, can generate high-quality distractors that are suitable for immediate use in an educational context.bitew2023distractor

### v2

@@ 1 Introduction
One of the most prominent advances in Artificial Intelligence recently is the development of large language models (LLMs), such as ChatGPT, BingChat, and Bard (developed by OpenAI, Microsoft, and Google, respectively). It can be said that large language models have been developing strongly in the past 2 years and creating a strong influence in the field of Generative AI (GenAI) Jahic et al. (2023).
LLMs are capable of solving a wide range of tasks, such as natural language understanding, text generation, and sentiment analysis in various domains Chang et al. (2023). Since the breakthroughs of AI and LLM, education, as a crucial role in shaping society through almost every single individual, might receive significant benefits from these new LLM-based initiatives. AI is transforming education, bridging its gaps, and promoting a more inclusive and productive learning environment by customizing learning experiences, automating administrative tasks, and providing real-time feedback Kamalov et al. (2023).
Advancements in AI and LLMs have fueled the development of many educational technology innovations that aim to automate the often time-consuming and laborious tasks of generating and analyzing textual content, including generating test questions. For the application of LLM in education, this technology holds great potential in this field at all levels, especially university level Yan et al. (2023). Therefore, this study conducts a research to comprehensively understand the capabilities of LLMs, especially ChatGPT when supporting lecturer to build question banks for learning modules in university setting. During the literature review, we witnessed that although the LLMs can provide quite a lot of support in teaching activities Xuan-Quy et al. (2023), the work of testing and evaluating learners, which takes up a lot of time and effort of lecturers, has not been studied in-depth by previous works. Therefore, in this paper, we decide to conduct research on the application of LLMs for generating question banks, with initial results at the Banking Academy of Vietnam. Specifically, the contributions of the paper are summarized as follows:
@@ 2 Related works
With the growing interest in LLM for education, many studies around the world focusing on this issue have been conducted and have produced some interesting results. A group of authors from the University of Minnesota Law School have conducted research on ChatGPT’s performance in answering a set of questions including 95 multiple-choice questions and 12 essay questions related to four undergraduate-level subjects in the law class. The results showed that ChatGPT studied and performed on the test at a level equivalent to a C+ grade for college students Jonathan et al. (2023).
Another study put ChatGPT to the test in completing the Dutch secondary school exam in the subject of English reading comprehension. This study concluded that ChatGPT achieved results with an average score of 7.3, comparable to the average score of all high school students in the country de Winter (2023).
In another research on evaluating the performance of LLMs when taking a high school level Biology test by the author namely Dao Xuan Quy, ChatGPT showed positive results with an accuracy of 71% for remembering level questions and 61.82% for understanding level questions. This result shows how effective ChatGPT is in capturing and clarifying concepts related to the required subject Xuan-Quy and Le (2023).
The use of LLMs in supporting the construction of questions to test and evaluate students’ knowledge should be encouraged. While exam-style questions are an important instructional tool for several reasons, manually creating questions is a time-consuming process that necessitates expertise, experience, and resources. This, in turn, impedes and inhibits the implementation of instructional activities (e.g., offering practice questions) and educational initiatives (e.g., adaptive testing) that need a huge pool of questions. Therefore, automatic question generating (AQG) approaches based on AI were established for research in both developed and developing nations Ghader et al. (2019).
@@
Selecting the optimal LLM-based solution . Nowadays, many advanced large language models can be widely applied in the field of education such as ChatGPT, BingAI, Gemini, etc. Most prominently, ChatGPT is an advanced general AI model based on a Generative Pretrained Transformer (GPT). This feature refers to the LLM being pre-trained with the available dataset and can generategrammatically and contextually human-like responses according to the user’s request through natural language processing. The current most popular versions of ChatGPT are GPT-3.5 and GPT-4, with the upgraded version GPT-4 can be used with a fee, which according to OpenAI performs much better with a 10 times larger pre-training datasets capacity Kalyan (2024).
The initial goal of this research is to evaluate various LLMs in order to find the one having performance and suitability for generating questions in the higher educational context. Each model has its own advantages and disadvantages, so our team has to compare consider which one can fulfill the both factors. As we prioritize the best ones meeting educational standards, ChatGPT Plus and Claude are two potential options. However, ChatGPT Plus has a relatively high fee (\$20/month) and Claude has a few limitations in its capabilities Lozić and Štular (2023).
@@
  • ChatGPT 3.5 has been proven through scientific research to be able to produce high-quality text, similar to text written by humansRay (2023).
  • ChatGPT 3.5 currently has 180 million users, with 100 million weekly active users and 1.6 billion website visit times Meer (2024). This number shows that ChatGPT 3.5 is sufficiently popular and trustworthy.
@@ 3.2 The investigation of appropriate Prompt patterns
Through research on the meaning of elements as well as the operation of available Prompt templates such as RTF (Role, Task, Format), RISE (Role, Input, Steps, Expectation), RTCF (Role, Task, Context, Format), RTCEF (Role, Task, Context, Example, Format). We have filtered, selected and separated into specific factors as shown in the Table 1 Čavojský et al. (2023).
@@ 3.5 The "Blind Test" method
To evaluate the quality of the questions more objectively, our research team conducted a survey using the “Blind test” method to collect answers from students to lecturers. Based on the well-known Turing Test, the "Blind test" aims to get empirical insight into people’s ability to discriminate between artificial and human content Köbis and Mossink (2021).
@@ 3.6.3 Specific analysis
Our data shows that the 9 questions having the least percentage of students answering correctly have an average of only 32.5% answers correct. In particular, the survey showed that the multiple-choice section had the lowest rate of correct answers with only about 30%. This may stem from the fact that human lecturers often create short, concise multiple-choice questions with the goal of simply testing whether students have grasped the knowledge surrounding the concept being asked. In fact, LLM like ChatGPT has the ability to grasp the subject and create related questions to support learnersXuan-Quy and Le (2023). Therefore, this tool is capable of creating MCQs properly along with the lecturers in an even faster way.
@@ D. Money loses value
Subjectively stated, a broad and general list of MCQs surveyed are less difficult to be guessed than real-life scenario exercises, or calculative ones that we had to give ChatGPT some examples to train it. Furthermore, these models can generate plausible distractors, which are incorrect answers that are designed to be similar to the correct answer, making it more challenging for humans to discern the difference. According to the study of Bitew et al. (2023), ChatGPT, when guided by question items and in-context examples, can generate high-quality distractors that are suitable for immediate use in an educational context.Bitew et al. (2023)